OpenStudy (anonymous):

In Equation such as 6x-y=29 and then i have x=y-9 i sub in y-9 6(y-9)-y=-29 solve for y which then equals 6y-54-y=-29 5y=25 ( how does it get their from 6y-54-y=-29?)

OpenStudy (anonymous):

and why

OpenStudy (gw2011):

In the original equation you show 6x-y=29 and int the other equation you show 6y-54-y=-29. I assume it should be -29 and not 29. 6y-54-y=-29 Add 54 to both sides of the equation in order to group like terms: 6y-54-y+54=-29+54 6y-y=25 5y=25 y=5